# Mirage: Evaluating and Explaining Inductive Reasoning Process in Language Models

### v1

@@ Results
We demonstrate the experimental results on GPT-4o in Table 3 and report other results in Appendix B.2. We can conclude that: (1) Advanced methods provide limited improvement to the model’s inductive reasoning ability or may even have negative effects. For both tasks, directly answering with few-shot settings can consistently achieve the highest or second-highest accuracy in most cases. After applying prompting methods like CoT, the model’s accuracy decreases by up to 18% and 22% on two tasks, respectively. It indicates that the key to optimizing inductive reasoning does not lie in refining the intermediate inductive process (as CoT-like methods do). (2) The model’s disregard for abstract rules during inductive reasoning is method-agnostic. Although some methods use instructions to guide the model to focus more on the previously induced rules during reasoning (e.g. ID, SR, HR), there remains a significant gap in the model’s inductive and deductive performance. For example, in the case of SR, the model’s deductive accuracy outperforms its inductive accuracy by an average of 16%.
@@ 3.3 Impact of Increasing Fact Size
Here, $\mathcal{M}(\cdot|I),\mathcal{M}(\cdot|D)$ are the model’s outputs in inductive and deductive tasks. We set $D$ = 5 and vary $N$, then analyze the distribution of these two thresholds across 100 samples, reporting the results in Figure 4.
@@
Figure 4: Performance on deductive tasks under different scenarios of observed and test facts.
@@ Results
Based on the result, we further prove that LLM’s deduction does not rely on an inductive rule. From both of the two figures, we can observe that most points are distributed in the upper left region of the line $x=y$, indicating that for the vast majority of cases, DCT is smaller than ICT. Therefore, the fact numbers $N$ does not affect the conclusion we stated earlier. LLM requires fewer facts to successfully perform a deductive task compared to correct induction.
@@
From the results shown in Figure 4, we can get that: (1) LLMs lack transferability in inductive reasoning. Across different cases, the highest performance occurs when the scenarios of the observed and test facts are consistent (i.e., the diagonal from the top left to the bottom right in the figure).(2) The inductive reasoning process of the LLM is form-related. Compared to the transfer between LT and RP (or CG and RP), the transfer between LT and CG demonstrates better performance. We infer that this is because the forms of LT and CG are more similar (see Figure 2). Based on the above two observations, we further confirm that LLMs do not rely on abstract rules when performing inductive reasoning. So, what is the underlying mechanism behind it? In the following section, we focus on addressing this question.
@@ 4.2 Neighbor Facts in Inductive Reasoning
Before conducting the experiments, we first define some key concepts in our work: the distance $d$ and neighborhood $\mathcal{N}$. In our setup, the components at corresponding positions in the vectors follow the same transformation rules, while non-corresponding components may undergo different transformations (see Equation 4). Hence, we consider using the distance based on the corresponding components: Chebyshev distance. [^6] We demonstrate through experiments that this distance is more suitable for constructing neighborhoods compared to other distances. For details, see Appendix C.2. Given observed fact ${\mathbb{X}}_{i}$ = $({\bm{x}}_{i},{\bm{y}}_{i})$ and test input ${\bm{x}}_{t}$, we have:
@@
According above definitions, we can divide an observed fact ${\mathbb{X}}_{k}$ into three categories based on the distance between ${\bm{x}}$ and the test input ${\bm{x}}_{t}$: (1) In-neighborhood Fact (IF): If ${\mathbb{X}}_{k}\in\mathcal{N}({\bm{x}}_{t},\epsilon)$, we call ${\mathbb{X}}_{k}$ is a in-neighborhood fact. (2) Cross-neighborhood Fact (CF): If ${\mathbb{X}}_{k}\notin\mathcal{N}({\bm{x}}_{t},\epsilon)$, but $\exists i\in[1,D],\ \text{s.t.}\ |{x}_{i}-{x}_{ti}|\leq\epsilon$, we consider it a suboptimal neighbor fact because some of its components can still contribute to the model’s inductive reasoning process. In this case, we call ${\mathbb{X}}_{k}$ is a cross-neighborhood fact. (3) Out-neighborhood Fact (OF): If $\forall i\in[1,D],\ |{x}_{i}-{x}_{ti}|>\epsilon$, we call ${\mathbb{X}}_{k}$ is an out-neighborhood fact. By substitution, we can make the fact set ${\mathbb{X}}$ contain only one type of the fact, while keeping the size $N$ fixed. After constructing different fact sets, we compare the model’s performance on deductive tasks under these settings. Besides, we use the performance under the default fact set as the baseline, where all facts are randomly sampled without any constraints.
@@ Results
We report the experimental results in Figure 5. It demonstrates that: (1) LLM’s inductive reasoning is neighbor-based. By comparing the three settings, we find that observed facts closer to the test case result in better performance (IF $>$ CF $>$ OF) across all of the models. Besides, compared to the baseline (the dashed line in figures), the accuracy significantly drops after we remove all the neighbor cases in ${\mathbb{X}}$ (i.e. OF). These phenomena indicate that the model heavily relies on neighbor facts during reasoning. (2) LLMs have a strong ability to capture neighboring patterns. When we set the neighborhood radius $\epsilon$ to 4, both IF and CF still contribute to high reasoning performances for the model. Besides, OF continues to show a significant decline (compared to $\epsilon$ = 3). These observations indicate that LLMs can still learn similar patterns even when the observed facts are relatively distant.
@@ 4.3 Universality of Neighbor-based Reasoning
We consider whether LLM’s inductive reasoning universally relies on neighbor cases, hence, we set $\epsilon$ to 1 and repeat the experiment under more different settings, where the baseline is the same as we set in §$\S$ 4.2. The results on GPT-4o are reported in Table 4, from which we can prove that the neighbor-based paradigm is universal in LLMs’ inductive reasoning process. Across different scenarios and fact numbers, IF consistently gets the highest accuracy, while OF gets the lowest accuracy. The reliance of LLMs’ inductive reasoning on neighbor facts is independent of the specific task scenarios, models, or fact numbers. For more results on other models, you can see Appendix C.3.
@@ 4.4 Effective Scope of Neighbor-based Reasoning
We have proved that neighbor examples in observed facts can significantly affect the model’s performance on the test case. However, what is the effective scope of it? Is it only effective on a single test example or does it generate an implicit rule that affects more examples? To answer the question, we first make three assumptions about its possible scope and show them in Figure 6. For individual scope, the model can only answer the test case ${\bm{x}}_{t}$ (e.g. [3,4,7] in the figure), for all other cases, the accuracy of the prediction is very low. For localized scope, the model can also successfully answer cases close to ${\bm{x}}_{t}$ (i.e. the neighbor facts of ${\bm{x}}_{t}$). For global scope, the model can answer all cases with high accuracy.
@@
In our experiment, we sample $n$ test cases ${\mathbb{X}}_{t}$ (here $n$ = 5) in each deductive task $\tau$ and define the accuracy $a_{\tau}$ for this particular task as:
@@
Here ${\mathbb{X}}_{\tau}$ is the observed fact set of the task. Let $T$ denote the set of all deductive tasks (we set $|T|$ = 100), we define deductive density $I_{d}$ as:
@@ 6 Related Work
Existing studies on evaluating LLM’s inductive reasoning capabilities mainly use only a single task. On one hand, some works assess the model’s rule induction ability by evaluating the accuracy on unseen examples (Moskvichev et al., 2023; Tang et al., 2023; Gendron et al., 2023; Xu et al., 2024b; Qiu et al., 2024). However, since the model’s deduction does not always rely on inducing the correct rule, this indirect evaluation method can introduce some inaccuracies. On the other hand, some studies directly evaluate the correctness of the generated rules to assess inductive reasoning ability (Shao et al., 2024; Cheng et al., 2024; Yang et al., 2024b). These studies lack evaluation on test examples, making it difficult to confirm the model’s mastery of the inductive rules. Our work evaluates both aspects, providing a comprehensive analysis of the model’s inductive reasoning process.
@@ 7 Conclusion
In this paper, we focus on evaluating and explaining the inductive reasoning process of LLMs. First, we construct a dataset Mirage, which provides both inductive and deductive evaluation tasks, with the flexibility to generate test examples in any distribution, different difficulties, and various forms. Based on it, we prove that LLM is a poor rule-based reasoner, it does not need to rely on inductive rules when performing deductive tasks. Compared to correct induction, the model can perform successful deduction with fewer observations, and this deduction is closely related to the form of the input. Furthermore, we identify a key paradigm of LLM inductive reasoning: neighbor-based reasoning. The model tends to leverage observed facts that are close to the test examples in feature space for inductive reasoning. Through it, the model can achieve strong inductive reasoning capabilities within a localized scope and apply this ability to make inferences on unseen examples.
@@ A.3 Evaluation Templates for Different Tasks and Scenarios
In Tables 13 and 14, we report the evaluation prompts used to evaluate the models in our work. In Table 15, we provide the templates used for constructing different scenarios in RP.
@@ B.2 More Experiments on other Models
Ind Ded ($\Delta$) Ind Ded ($\Delta$) Ind Ded ($\Delta$) Ind Ded ($\Delta$)
@@
Ind Ded ($\Delta$) Ind Ded ($\Delta$) Ind Ded ($\Delta$) Ind Ded ($\Delta$)
@@ C.2 Comparison with Other Distance Metric
where we set $p=3$. We can generate three distinct new neighborhoods $\mathcal{N}({\bm{x}}_{t},\epsilon)$ by incorporating these distances into Equation 9, thereby constructing three new kinds of OF. Therefore, we compare the model’s performance on deductive tasks when using only these different OFs, and the results are shown in Figure 8. From the figure, we can see that our neighborhood construction outperforms those constructed using other distance metrics. The deductive performance of the other three OFs across different radii is similar to the baseline, indicating that removing neighbor facts constructed using these methods does not influence the model’s inductive reasoning ability. In contrast, our constructed OF leads to a significant decline in accuracy, proving the validity of our neighborhood construction.
@@ C.3 More Experiments on other Models
We repeat the experiments in §$\S$ 4.3 on Llama2-13B, Claude-3.5 and Llama3-8B. The results are shown in Table 9, 10, 11. Besides, we also repeat the experiments in §$\S$ 4.4 on Claude-3.5 and report the results in Figure 9. The results of all these additional experiments are consistent with those in the main text.
@@ C.4 Supplementary Experiment for Main Experiment
We observe that, in the experiment of §$\S$ 4.2, though the performance of OF significantly decreases compared to the baseline, some models are still able to maintain around 40% accuracy, even with only distant observed facts. We infer that models are likely to conduct rule-based reasoning in these cases. Hence, we design an extra experiment for supplementary. In it, we prompt LLMs to induct rules and finish deductive tasks (i.e. ID in §$\S$3.2) on these cases in Table 12. From the table, we can observe that the model’s deductive accuracy using the rule exceeds 70% when there are fewer neighbor facts in the context. This demonstrates that the model tends to rely more on rule-based induction if there is less neighbor-based matching.
@@
Table 13: Prompts for inductive tasks ($D$ = 3).
@@
Table 14: Prompts for deductive tasks ($D$ = 3).

### v2

@@ Results
We illustrate the main results on GPT-4o in Table 3 (more results in Appendix B.2), from which we can conclude that: (1) Advanced methods provide limited improvement to the model’s inductive reasoning ability or may even have negative effects. For both tasks, directly answering with few-shot settings can consistently achieve the highest or second-highest accuracy in most cases. After applying methods like CoT, the model’s accuracy decreases by up to 18% and 22% on two tasks, respectively. It indicates that the key to optimizing inductive reasoning does not lie in refining the intermediate inductive process (as CoT-like methods do). (2) The model’s disregard for abstract rules during inductive reasoning is method-agnostic. Although some methods use instructions to guide the model to focus more on the induced rules during reasoning (e.g. ID, SR, HR), there remains a significant gap in the model’s RI and EI performance. For example, in the case of SR, the model’s example inference accuracy outperforms its rule induction accuracy by an average of 16%.
@@ 3.3 Impact of Increasing Fact Size
Here, $\mathcal{M}(\cdot|I),\mathcal{M}(\cdot|D)$ are the model’s outputs in RI and EI tasks. We set $D$ = 5 and vary $N$, then analyze the distribution of these two thresholds across 100 samples, reporting the results in Figure 4.
@@
Figure 4: Performance on EI tasks under different scenarios of observed and test facts.
@@ Results
Based on the result, we further demonstrate that LLM’s deduction does not rely on an inductive rule. From both of the two figures, we can observe that most points are distributed in the upper left region of the line $x=y$, indicating that for the vast majority of cases, DCT is smaller than ICT. Therefore, the fact numbers $N$ does not affect the conclusion we stated earlier. LLM requires fewer facts to successfully perform an example inference task compared to correct induction.
@@
From the results shown in Figure 4, we can get that: (1) LLMs lack transferability in inductive reasoning. Across different cases, the highest performance occurs when the scenarios of the observed and test facts are consistent (i.e., the diagonal from the top left to the bottom right in the figure).(2) The inductive reasoning process of the LLM is form-related. Compared to the transfer between LT and RP (or CG and RP), the transfer between LT and CG demonstrates better performance. We infer that this is because the forms of LT and CG are more similar (see Figure 2). In addition, we also design experiments under the fine-tuning paradigm to compare the model’s transferability (presented in Appendix B.3), and the results remain consistent. Based on the above two observations, we further confirm that LLMs do not rely on abstract rules when performing inductive reasoning. So, what is the underlying mechanism behind it? In the following section, we focus on addressing this question.
@@ 4.2 Neighbor Facts in Inductive Reasoning
Before conducting the experiments, we first define some key concepts in our work: the distance $d$ and neighborhood $\mathcal{N}$. In our setup, the components at corresponding positions in the vectors follow the same transformation rules, while non-corresponding components may undergo different transformations (see Equation 4). Hence, we consider using the distance based on the corresponding components: Chebyshev distance (further discussion in Appendix C.2). Given observed fact ${\mathbb{X}}_{i}$ = $({\bm{x}}_{i},{\bm{y}}_{i})$ and test input ${\bm{x}}_{t}$, we have:
@@
According above definitions, we can divide an observed fact ${\mathbb{X}}_{k}$ into three categories based on the distance between ${\bm{x}}$ and the test input ${\bm{x}}_{t}$: (1) In-neighborhood Fact (IF): If ${\mathbb{X}}_{k}\in\mathcal{N}({\bm{x}}_{t},\epsilon)$, we call ${\mathbb{X}}_{k}$ is a in-neighborhood fact. (2) Cross-neighborhood Fact (CF): If ${\mathbb{X}}_{k}\notin\mathcal{N}({\bm{x}}_{t},\epsilon)$, but $\exists i\in[1,D],\ \text{s.t.}\ |{x}_{i}-{x}_{ti}|\leq\epsilon$, we consider it a suboptimal neighbor fact because some of its components can still contribute to the model’s inductive reasoning process. In this case, we call ${\mathbb{X}}_{k}$ is a cross-neighborhood fact. (3) Out-neighborhood Fact (OF): If $\forall i\in[1,D],\ |{x}_{i}-{x}_{ti}|>\epsilon$, we call ${\mathbb{X}}_{k}$ is an out-neighborhood fact. By generating examples based on these definitions, we can make the fact set ${\mathbb{X}}$ contain only one type of fact. After constructing different fact sets, we compare the model’s performance on EI tasks under these settings. Besides, we use the performance under the default fact set as the baseline, where all facts are randomly sampled without any constraints.
@@ Results
We report the results in Figure 5. It demonstrates that: (1) LLM’s inductive reasoning is neighbor-based. By comparing three settings, we find that observed facts closer to the test case result in better performance (IF $>$ CF $>$ OF) across all models. Besides, compared to the baseline (the dashed line in figures), the accuracy significantly drops after removing the neighbor cases in ${\mathbb{X}}$ (i.e. OF). These phenomena indicate that the model heavily relies on neighbor facts during reasoning. (2) LLMs have a strong ability to capture neighboring patterns. When we set the neighborhood radius $\epsilon$ to 4, both IF and CF still contribute to high accuracy for the model. Besides, OF continues to show a significant decline (compared to $\epsilon$ = 3). These observations indicate that LLMs can still learn similar patterns even when the observed facts are relatively distant.
@@ 4.3 Universality of Neighbor-based Reasoning
We consider whether LLM’s inductive reasoning universally relies on neighbor cases, hence, we set $\epsilon$ to 1 and repeat the experiment under different settings, where the baseline is the same as §$\S$ 4.2. The results on GPT-4o are reported in Table 4, which demonstrates that the neighbor-based paradigm is universal in LLMs’ inductive reasoning process. Across different scenarios and fact numbers, IF consistently gets the highest accuracy, while OF gets the lowest accuracy. The reliance of LLMs’ inductive reasoning on neighbor facts is independent of the specific task scenarios, models, or fact numbers. We present more results in Appendix C.3.
@@ 4.4 Effective Scope of Neighbor-based Reasoning
We have demonstrated that neighbor examples in observed facts significantly affect the model’s performance on the test case. However, what is the effective scope of it? Is it only pattern matching on a single test example or reasoning an implicit rule that affects more examples? To answer the question, we first make three assumptions about its possible scope and show them in Figure 6. For individual scope, the model can only answer the test case ${\bm{x}}_{t}$ (e.g. [3,4,7] in the figure), for all other cases, the accuracy of the prediction is very low. For localized scope, the model can also answer cases close to ${\bm{x}}_{t}$ (i.e. the neighbor facts of ${\bm{x}}_{t}$). For global scope, the model can answer all cases with high accuracy.
@@
In our experiment, we sample $n$ test cases ${\mathbb{X}}_{t}$ (here $n$ = 5) in each EI task $\tau$ and define the accuracy $a_{\tau}$ for this particular task as:
@@
Here ${\mathbb{X}}_{\tau}$ is the observed fact set of the task. Let $T$ denote the set of all EI tasks (we set $|T|$ = 100), we define deductive density $I_{d}$ as:
@@ 6 Related Work
Existing studies on evaluating LLM’s inductive reasoning capabilities mainly use only a single task. On one hand, some works assess the model’s rule induction ability by evaluating the inference accuracy on unseen examples (Moskvichev et al., 2023; Tang et al., 2023; Gendron et al., 2023; Xu et al., 2024b; Qiu et al., 2024). However, since the model’s deduction does not always rely on inducing the correct rule, this indirect evaluation method can introduce some inaccuracies. On the other hand, some studies directly evaluate the correctness of the generated rules to assess inductive reasoning ability (Shao et al., 2024; Cheng et al., 2024; Yang et al., 2024b; Wang et al., 2025). These studies lack evaluation on test examples, making it difficult to confirm the model’s mastery of the inductive rules. Our work evaluates both aspects, providing a comprehensive analysis of the model’s inductive reasoning process.
@@ 7 Conclusion
In this paper, we focus on evaluating and explaining the inductive reasoning process of LLMs. First, we construct a dataset Mirage, which provides both inductive and deductive evaluation tasks, with the flexibility to generate test examples in any distribution, different difficulties, and various forms. Based on it, we demonstrated that LLM is a poor rule-based reasoner, it does not need to rely on inductive rules when performing inductive reasoning. Compared to correct induction, the model can perform successful deduction with fewer observations, and this deduction is closely related to the form of the input. Furthermore, we identify a key paradigm of LLM inductive reasoning: neighbor-based reasoning. The model tends to leverage observed facts that are close to the test examples in feature space for inductive reasoning. Through it, the model can achieve strong inductive reasoning capabilities within a localized scope and apply this ability to make inferences on unseen examples.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No.U24A20335, No. 62176257, No. 62406321). This work is also supported by Beijing Natural Science Foundation (L243006). This work is also supported by the Youth Innovation Promotion Association CAS and the China Postdoctoral Science Foundation under Grant Number 2024M753500.
@@ A.3 Evaluation Templates for Different Tasks and Scenarios
In Tables 14 and 15, we report the evaluation prompts used to evaluate the models in our work. In Table 16, we provide the templates used for constructing different scenarios in RP.
@@ B.2 More Experiments on other Models
RI EI ($\Delta$) RI EI ($\Delta$) RI EI ($\Delta$) RI EI ($\Delta$)
@@
RI EI ($\Delta$) RI EI ($\Delta$) RI EI ($\Delta$) RI EI ($\Delta$)
@@
B.3 Experiments on Fine-tuning Method
In the main text, we primarily use in-context learning to evaluate the performance of various models. Here, we supplement the evaluation with the performance of fine-tuned models on Mirage. Specifically, we use Meta-Llama-3-8B-Instruct as the backbone, setting N = 5 and D = 5, and train the model on 8,000 samples. For the training parameters, we set the learning rate to 0.0001, the batch size to 1, and the number of epochs to 10. Additionally, LoRA is employed to train people on different types of tasks. The performances on 100 test examples are presented in Table 9. It demonstrates that fine-tuning can effectively enhance the model’s inductive reasoning capabilities. Compared to the 5-shot ICL, the model performs better on both reasoning tasks, even when not trained on tasks of the same format. However, consistent with the conclusion in §$\S$ 3.4, the results of the training do not exhibit good formal generalization. The model tends to perform relatively poorly on test sets with significantly different forms from the training set (e.g., LT and RP).
Method LT RP CG ST
RI EI RI EI RI EI RI EI
5-shot ICL 0.31 0.27 0.04 0.17 0.16 0.28 0.21 0.32
LT Training 0.89 0.82 0.22 0.24 0.34 0.80 0.26 0.38
RP Training 0.51 0.44 0.78 0.74 0.42 0.69 0.24 0.35
CG Training 0.76 0.75 0.22 0.25 0.86 0.80 0.26 0.37
ST Training 0.51 0.52 0.19 0.21 0.33 0.61 0.50 0.35
Table 9: Performance of the fine-tuned model on Mirage. The best results in each column are highlighted in bold.
@@ C.2 Comparison with Other Distance Metric
where we set $p=3$. We can generate three distinct new neighborhoods $\mathcal{N}({\bm{x}}_{t},\epsilon)$ by incorporating these distances into Equation 9, thereby constructing three new kinds of OF. Therefore, we compare the model’s performance on EI tasks when using only these different OFs, and the results are shown in Figure 8. From the figure, we can see that our neighborhood construction outperforms those constructed using other distance metrics. The EI performance of the other three OFs across different radii is similar to the baseline, indicating that removing neighbor facts constructed using these methods does not influence the model’s inductive reasoning ability. In contrast, our constructed OF leads to a significant decline in accuracy, proving the validity of our neighborhood construction.
@@ C.3 More Experiments on other Models
We repeat the experiments in §$\S$ 4.3 on Llama2-13B, Claude-3.5 and Llama3-8B. The results are shown in Table 10, 11, 12. Besides, we also repeat the experiments in §$\S$ 4.4 on Claude-3.5 and report the results in Figure 9. The results of all these additional experiments are consistent with those in the main text.
@@ C.4 Supplementary Experiment for Main Experiment
We observe that, in the experiment of §$\S$ 4.2, though the performance of OF significantly decreases compared to the baseline, some models are still able to maintain around 40% accuracy, even with only distant observed facts. We infer that models are likely to conduct rule-based reasoning in these cases. Hence, we design an extra experiment for supplementary. In it, we prompt LLMs to induct rules and finish example inference tasks (i.e. ID in §$\S$3.2) on these cases in Table 13. From the table, we can observe that the model’s deductive accuracy using the rule exceeds 70% when there are fewer neighbor facts in the context. This demonstrates that the model tends to rely more on rule-based induction if there is less neighbor-based matching.
@@
Table 14: Prompts for rule induction tasks ($D$ = 3).
@@
Table 15: Prompts for example inference tasks ($D$ = 3).